ou deposit P200 in an account earning 5% compound interest for 3 years. Find the future value and the interest earned for each of the following compounding frequencies. Use the Bankers' Rule for daily compounding.

Frequency Future Value Interest Earned
Annually:
Semiannually:
Quarterly:
Monthly:
Daily:

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

annually is P=200(1+0.05/1)^3=P231.53
semi is P=200(1+(.05/2))^6, less to compound but more times compounded,=P231.94
quarterly will be slightly more or P=200(1+(.05/4))^12=P232.15
monthly is P=200*(1+(.05/12))^36=P232.29
Round everything at the end.
daily use 360 so P=(1+(.05/360))^1080=P232.36

Answer by ikleyn(52776)

.

To see many other similar  (and different)  solved problems on compounded interest accounts,  look into the lesson
    - Compounded interest percentage problems
in this site.


Learn the subject from there.